# Image2Net: Datasets, Benchmark and Hybrid Framework to Convert Analog Circuit Diagrams into Netlists

### v1

@@ Abstract
Large Language Model (LLM) exhibits great potential in designing of analog integrated circuits (IC) because of its excellence in abstraction and generalization for knowledge. However, further development of LLM-based analog ICs heavily relies on textual description of analog ICs, while existing analog ICs are mostly illustrated in image-based circuit diagrams rather than text-based netlists. Converting circuit diagrams to netlists help LLMs to enrich the knowledge of analog IC. Nevertheless, previously proposed conversion frameworks face challenges in further application because of limited support of image styles and circuit elements. Up to now, it still remains a challenging task to effectively convert complex circuit diagrams into netlists. To this end, this paper constructs and opensources a new dataset with rich styles of circuit diagrams as well as balanced distribution of simple and complex analog ICs. And a hybrid framework, named Image2Net, is proposed for practical conversion from circuit diagrams to netlists. The netlist edit distance (NED) is also introduced to precisely assess the difference between the converted netlists and ground truth. Based on our benchmark, Image2Net achieves 80.77% successful rate, which is 34.62%-45.19% higher than previous works. Specifically, the proposed work shows 0.116 averaged NED, which is 62.1%-69.6% lower than state-of-the-arts.
@@
Index Terms:
Analog Circuit Design, Dataset, EDA, Netlist, Object Detection, Valuation Method
@@ II-B Evaluation Metric: NED
Successful $\text{Successful Rate}=\frac{N_{success}}{N_{total}}$
@@ IV-A NED Result
To evaluate our algorithm, we generated the netlists using our algorithm on our public test dataset at II-A4 and calculated the NED in comparison with the golden data. If the result is all correct, the NED should be 00. After providing each case with sufficient running time, according to our evaluation method in chapter II-B, we calculated the NED for each case. Our solution ultimately achieved an accuracy rate of 80.77%, and the $\overline{NED}$ is 0.659. Score distribution is presented in Fig.7. According to that the calculation of GED is a NP-Hard problem, we use the algorithm mentioned in II-B and calculated for 2 hours for each case.
@@
Method Successful Rate $\uparrow$ $\overline{NED}$downarrow
@@ Appendix B Complete definition of our device and crossing types
MOS NMOS nmos Gate, Source, Drain, (Body)
@@
BJT NPN npn Base, Collector, Emmitter
@@
DIDO_Amp dido_amp InP, InN, OutP, OutN
@@ E-A Class-Specific Performance
Accuracy=True PositivesTrue Positives+False $\text{Accuracy}=\frac{\text{True Positives}}{\text{True Positives}+\text{FalsePositives}}$
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$
@@ E-B Target Detection Performance
[…]
@@ E-D Crossing Identification
[…]

### v2

@@ Abstract
Large Language Model (LLM) exhibits great potential in designing of analog integrated circuits (IC) because of its excellence in abstraction and generalization for knowledge. However, further development of LLM-based analog ICs heavily relies on textual description of analog ICs, while existing analog ICs are mostly illustrated in image-based circuit diagrams rather than text-based netlists. Converting circuit diagrams to netlists help LLMs to enrich the knowledge of analog IC. Nevertheless, previously proposed conversion frameworks face challenges in further application because of limited support of image styles and circuit elements. Up to now, it still remains a challenging task to effectively convert complex circuit diagrams into netlists. To this end, this paper constructs and opensources a new dataset with rich styles of circuit diagrams as well as balanced distribution of simple and complex analog ICs. And a hybrid framework, named Image2Net, is proposed for practical conversion from circuit diagrams to netlists. The netlist edit distance (NED) is also introduced to precisely assess the difference between the converted netlists and ground truth. Based on our benchmark, Image2Net achieves 80.77% successful rate, which is 34.62%-45.19% higher than previous works. Specifically, the proposed work shows 0.116 averaged NED, which is 62.1%-69.6% lower than state-of-the-arts. Our datasets and benchmark are available at https://github.com/LAD021/ci2n_datasets.
@@ II-B Evaluation Metric: NED
$\text{Successful Rate}=\frac{N_{success}}{N_{total}}$
@@ IV-A NED Result
To evaluate our algorithm, we generated the netlists using our algorithm on our public test dataset at II-A4 and calculated the NED in comparison with the golden data. If the result is all correct, the NED should be 0. After providing each case with sufficient running time, according to our evaluation method in chapter II-B, we calculated the NED for each case. Our solution ultimately achieved an accuracy rate of 80.77%, and the $\overline{NED}$ is 0.659. Score distribution is presented in Fig.7. According to that the calculation of GED is a NP-Hard problem, we use the algorithm mentioned in II-B and calculated for 2 hours for each case.
@@
Method Successful Rate $\uparrow$ $\overline{NED}$ $\downarrow$
@@ Appendix B Complete definition of our device and crossing types
MOS NMOS nmos
Gate,
Source,
Drain,
@@
BJT NPN npn
Base,
Collector,
Emmitter
@@
DIDO_Amp dido_amp
InP, InN,
OutP, OutN
@@ E-A Class-Specific Performance
$\text{Accuracy}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$
$\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$
@@ E-B Target Detection Performance
$\text{mAP@50-95}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{10}\sum_{j=0.5}^{0.95}\Delta_{j}\cdot\text{AP}_{\text{class}_{i}}(j)\right)$
@@ E-D Crossing Identification
$\text{mAP@inside}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}_{\text{class}_{i}}(acip)$